a pollen sample weighs 12 grams
and contains 1:5 × 10 to the power nine pollen grains.
calculate the weight of one pollen grain in grams.
give your answer in scientific notation.

Found 3 solutions by MathLover1, math_tutor2020, ikleyn:
Answer by MathLover1(20849)
Answer by math_tutor2020(3816)


12 = 1.2 * 10^1
1.5*10^9 = 1.5 billion = 1,500,000,000
There are 8 zeros after the "1,5" portion.

x = weight, in grams, of 1 pollen grain
(1.5*10^9)x = weight of 1.5*10^9 grains = weight of 1.5 billion grains

(1.5*10^9)x = 12 grams
(1.5*10^9)x = (1.2*10^1) grams
(1.5*10^9)x = 1.2*10^1
x = (1.2*10^1)/(1.5*10^9)
x = (1.2/1.5)*10^(1-9)
x = 0.8*10^(-8)
x = (8.0*10^(-1))*10^(-8)
x = 8.0*( 10^(-1)*10^(-8) )
x = 8.0*10^( -1+(-8) )
x = 8.0*10^( -9 )
This is 8 billionths of a gram.
Be sure to not mix up the place value "billions" with "billionths".

How can we interpret such a small number?
If the single pollen grain was 1 billionth of a gram, then we'd need 1 billion grains to get 1 gram collectively.
It's similar to how we need 1000 millimeter pieces to make up a full meter.

Side note: many calculators offer the ability to switch to scientific notation mode.

-------------------------------------------------------------
-------------------------------------------------------------

Answer: 8.0 * 10^(-9) grams

Answer by ikleyn(52780)

.

As  I  (and other tutors)  see this format in your post  " 1:5 × 10 to the power nine ",  it seems strange.

A normal writing is  (probably)  " 1.5 × 10 to the power nine ".  It is how me  (and other tutors)  read it.


        So,  we all read it differently from what we see.
                 It is just bordering with a crime.


        (By the way, it is you who pushes us to this crime)


But what your writing actually means,  we do not know and can not know,  since you use non-standard format.

            (It is the same as if you use chines' hieroglyphs)


                                We can only guess.


In  Math,  every word,  every symbol and every punctuation sign does matter.


                    And if some post contains non-standard formatting,
         it should be rejected,  strictly saying,  and reported up to the instance.


So,  my question is:  from which source did you re-write your number
and what,  ACTUALLY,  is written in that source ?